OpenStudy (anonymous):

if 2x/x^2+1 = 2/x+2, what is the value of x? a) -1/4 b)1/4 c)1/2 d) 0 e)2

OpenStudy (anonymous):

Are there supposed to be brackets anywhere?

OpenStudy (anonymous):

no brackets

OpenStudy (anonymous):

Is it \[2x/(x^2+1) = 2/(x+2)\]?? or not?

OpenStudy (anonymous):

i guess you can look at it that way

OpenStudy (anonymous):

Well without the brackets, x can be any number except zero.

OpenStudy (anonymous):

2x/(x2+1)=2/(x+2) x/(x2+1)=1/(x+2) (x2+1)/x=(x+2) (x2+1)=x(x+2) x2+1=x2+2x 2x=1 x=1/2